The question is: Almonds worth $8.75/lb mixed with peanuts worth 4.75/lb take how many pounds of each to make an 80lb mixture worth $6.50/lb?
I have tried solving by addition with 8.75a + 4.75p = 80(6.50) or $520 by mutiplying a+b=80 by -4.75 to eliminate the peanuts.The result I get is 4.75a=$520-$380 or $140. $140/4.75 is 29.47 for almonds which is wrong. I have the correct answer and my calculation doesn't match the answer and I don't know what I'm doing wrong. I need to know the right way to get to 45lbs of almonds and 35 lbs of peanuts.
This is not from a textbook but a final review.
Thank you very much

Answer by Edwin McCravy(20086)


 8.75a + 4.75p =  520
-4.75a - 4.75p = -380
----------------------
    4a         =  140


Your error was that you had 4.75a where you should
have only had 4a. Note that 8.75-4.75 = 4, not 4.75.


 Answer a = 35, p = 45

Edwin